OpenStudy (anonymous):

For the matrices given in numbers 1-5, find the value of each determinant: |2 3 | |-1 4| Select one: a. 11 b. 4 c. 5 d. -22

OpenStudy (nurali):

D=ad-bc D=(2)(4)-(3)(-1) D=8+3 D=11

OpenStudy (cwrw238):

this is the way you do these